mass of acetic acid?
molar mass=60.052 g/mol
heat capacity=8.43kj/C
temp increase from 24.5 C to 36.8 C
deltarU for combustion =-874.2kj/mol
1
Expert's answer
2018-12-11T03:59:54-0500

Answer on Question #83705, Chemistry/ Organic Chemistry

mass of acetic acid?

molar mass=60.052 g/mol

heat capacity=8.43kj/C

temp increase from 24.5 C to 36.8 C

deltarU for combustion =-874.2kj/mol

Solution:

ΔErxn = qrxn = -qcal = -Ccal*Δt = -Ccal*(TF - TI) = - 8.43 kJ/oC *(36.8 oC - 24.5oC) = - 103.7 kJ

ΔUcomb = -874.2 kJ/mol

When 1 mole of acetic acid is combusted, 874.2 kJ of energy is absorbed by calorimeter

When x moles of acetic acid is conbusted, 103.7 KJ of energy is absorbed by calorimeter

Solve the proportion:

1/x = 874.2/103.7

x= 0.119

n(acetic acid) = 0.119 moles

m= M*n

m(acetic acid) = 0.119 moles * 60.052 g/mol =7.146 g

Answer: 7.146 g